OpenStudy (anonymous):

If sin theta=3/5 and theta is in quad 2 the exact form of csc (pi/2-theta) is...? Just want to make sure if this works??

OpenStudy (anonymous):

sinθ = 3/5 csc(π/2 - θ) = secθ = 1/cosθ = -4/5

OpenStudy (triciaal):